# FNOPE: Simulation-based inference on function spaces with Fourier Neural Operators

### v1

@@ Conclusion
Figure S2: Posterior samples Darcy flow experiment. Additional posterior samples for the Darcy flow experiment, for six distinct observations simulated from ground truth parameter $\theta_{i}$.
Figure S3: Posterior means for Darcy flow experiment. Posterior means (based on 100 samples) for the Darcy flow experiment, for six distinct observations simulated from ground truth parameter $\theta_{i}$.
Figure S4: Posterior standard deviation for Darcy flow experiment. Posterior standard deviations (based on 100 samples) for the Darcy flow experiment, for six distinct observations simulated from ground truth parameter $\theta_{i}$.
Figure S5: Mass balance rates of ice shelves. Inference of mass balance rates where the parameter discretization uses 500 gridpoints (observation discretization remains unchanged). Performance measures on test simulations and the real observation, where NPE (raw) refers to the method used in Moss et al. [39]. Results for FMPE (raw) omitted as this baseline was not able to always produce samples within the prior bounds across all test observations. See Fig. 6 for the results based on 50 grid points.

### v2

@@ Conclusion
S8.1 Additional Darcy results
Figure S6: Posterior samples Darcy flow experiment. Additional posterior samples for the Darcy flow experiment for six distinct observations simulated from ground truth parameter $\theta_{i}$.
Figure S7: Posterior means for Darcy flow experiment. Posterior means (based on 100 samples) for the Darcy flow experiment for six distinct observations simulated from ground truth parameter $\theta_{i}$.
Figure S8: Posterior standard deviation for Darcy flow experiment. Posterior standard deviations (based on 100 samples) for the Darcy flow experiment for six distinct observations simulated from ground truth parameter $\theta_{i}$.
S8.2 Comparison to invertible Fourier Neural Operator
As an additional comparison, we apply the work of long2025iFNO and train an invertible Fourier Neural Operator (iFNO) to solve the Darcy Flow inverse problem, using the same training data and simulation budgets. We train the iFNO with the same settings as described in long2025iFNO for the D-CURV experiment, summarized below. We observe that while iFNO achieves a comparable performance to FNOPE in terms of its predictive MSE, the posterior distribution is not calibrated as measured by the SBC EoD (Fig. S9a,b) and essentially collapsed to a point estimate. The SBC EoD value measured for iFNO (around 0.25) is consistent with this point estimate, because the one-dimensional marginal of a point mass distribution either always overestimates or underestimates the ground truth value. Hence, recalling the definition of SBC EoD (Sec. S2.2), the ground truth will have rank $r_{ij}=1$ or $r_{ij}=K_{\text{post}}+1$. Supposing that over different observations $x_{j}^{o}$, the point mass distribution is equally likely to underestimate or overestimate the ground truth, the cumulative distribution of the ranks will be given by $\text{CDF}_{i}(\alpha)=0.5$ for all significance levels $\alpha\in(0,1)$. Given the definition of the SBC EoD, a point mass distribution results in a SBC EoD value of
$\int_{0}^{1}|\text{CDF}_{i}(\alpha)-\alpha|d\alpha=0.25$ (7)
for each dimension $i$. This overconfidence is also reflected in the standard deviations of the estimated distributions (Fig. S9c), which show that iFNO essentially estimates a point mass for this task.
iFNO hyperparameters
We trained iFNO with 4 FNO blocks and 16 Fourier modes. The number of training epochs for the invertible Fourier blocks was set to 100, for the $\beta$-VAE 1000, and for joint training 100. The architecture of the $\beta$-VAE was the same as in long2025iFNO with rank 32, as well as the value $\beta=10^{-6}$. We used a minibatch size of $10$ for joint training and $20$ for the VAE and iFNO pretraining.
Figure S9: Darcy Flow, comparison of FNOPE to iFNO [long2025iFNO]. (a) MSE of posterior predictives to the ground truth observation (zoomed in relative to Fig. 5). (b) Simulation-based calibration Error of Diagonal (SBC EoD) for different training budgets. (c) Ground truth $\theta$, Posterior means, pixelwise error of means relative to ground truth, and posterior standard deviations. The color bars of the means match Fig. S7, and for errors and standard deviations they match Fig. S8.